# Quantum Wave Function Collapse for Procedural Content Generation

### v1

@@ Abstract
Quantum computers exhibit an inherent randomness, so it seems natural to consider them for procedural content generation. In this work, a quantum version of the famous (classical) wave function collapse algorithm is proposed. This quantum wave function collapse algorithm is based on the idea that a quantum circuit can be prepared in such a way that it acts as a special-purpose random generator for content of a desired form. The proposed method is presented theoretically and investigated experimentally on simulators and actual IBM Quantum devices.
@@
Originally proposed in Ref. [1], the WFC algorithm is a a non-backtracking, greedy constraint solving method [2, 3] that is able to generate complex patterns based on a set of input samples. It is known for its ability to create diverse and complex outputs that resemble the input samples while exhibiting novel combinations and variations. WFC employs two implementation strategies: the simple tiled model and the overlapping model, which share an identical algorithm core [4]. In the simple tiled model, tilesets are manually prescribed with predefined adjacency constraints, whereas the overlapping model automatically generates this information from a sample input.
@@
Despite its name, WFC is a purely classical algorithm. But even if WFC has nothing to do with quantum physics beyond the terminology, the question naturally arises as to whether quantum computers can be used to execute a genuine quantum version of the algorithm. The investigation of this question and a first draft for a corresponding algorithm are the subject of this manuscript. For this purpose, the terms classical wave function collapse (CWFC), quantum wave function collapse (QWFC), and hybrid quantum-classical wave function collapse (HWFC) are used In the following to clearly distinguish between classical, quantum and hybrid quantum-classical WFC.
Leveraging quantum computers for PCG has been considered before, mainly focusing on the quantum generalization of a blurring process [5, 6] and map generation using a quantum-enhanced decision making process [7]. Related projects have been carried out in connection with the development of quantum-inspired games [8].
The contribution of this paper is threefold:
@@
  • A probabilistic formulation of the simple tiled model of CWFC is presented as a foundation for a quantum version.
  • A QWFC method and a HWFC method are proposed.
  • The proposed methods are tested on simulators and actual IBM Quantum devices.
@@
The remaining manuscript is structured as follows. In the next section, a formal description of CWFC is provided in an appropriate probabilistic form, which enables the development of a QWFC method and a HWFC method in the subsequent section. In the section after, the proposed methods are demonstrated in practice. Finally, the paper ends with a conclusion and an outlook.
1 CLASSICAL METHODS
In this section, the formal description for a probabilistic iterative PCG (PIPCG) is provided and, subsequently, it is shown how it can be reduced to the common CWFC approach.
1.1 PIPCG
@@
where each segment $(i,v_{i})$ is defined by a unique identifier $i\in[1,N]$ and a value $v_{i}\in A_{i}$ from an alphabet $A_{i}\subset\mathbb{N}$. Here and in the following, the notation $[i,j]:=\{i,\dots,j\}\,\forall\,i,j\in\mathbb{N}$ is used. The set of all possible content instances $\mathcal{C}$ is determined by the content alphabets $\mathcal{A}:=\{A_{1},\dots,A_{N}\}$ for a given size $N$ with $|\mathcal{C}|=\prod_{i=1}^{N}A_{i}$. For example, if the content represents an image, its individual pixels can constitute the segments. Each pixel can attain one color from the available color space of the image (as its alphabet).
1.1.1 Iterative process.
@@
    Randomly choose a new segment identifier $s(k)\in[1,N]$ from a set of available identifiers.
@@
    Randomly choose a corresponding value $v_{s(k)}\in A_{s(k)}$ from a set of available values.
@@
By repeating these steps, the content instance is assembled in the sense of $C(0)\overset{k=1}{\rightarrow}C(1)\overset{k=2}{\rightarrow}\cdots\overset{k=N}{\rightarrow}C(N)$, where
$C(k):=\{(s(1),v_{s(1)}),\dots,(s(k),v_{s(k)})\}\in\mathcal{C}(k),$ (2)

### v2

# angle=0, scale=1, contents= This article has been accepted for publication in IEEE Computer Graphics and Applications. This is the author’s version which has not been fully edited and

@@ Abstract
Quantum computers exhibit an inherent randomness, so it seems natural to consider them for procedural content generation. In this work, a quantum version of the famous (classical) wave function collapse algorithm is proposed. This quantum wave function collapse algorithm is based on the idea that a quantum circuit can be prepared in such a way that it acts as a special-purpose random generator for content of a desired form. The proposed method is presented theoretically and investigated experimentally on simulators and IBM Quantum devices.
[…]
@@
The WFC algorithm [1] is a a non-backtracking, greedy constraint solving method [2, 3] that is able to generate complex patterns based on a set of input samples. It is known for its ability to create diverse and complex outputs that resemble the input samples while exhibiting novel combinations and variations. WFC employs two implementation strategies: the simple tiled model and the overlapping model, which share an identical algorithm core [4]. In the simple tiled model, tilesets are manually prescribed with predefined adjacency constraints, whereas the overlapping model automatically generates this information from a sample input.
@@
Despite its name, WFC is a purely classical algorithm. But even if WFC has nothing to do with quantum physics beyond the terminology, the question arises as to whether quantum computers can still be used to execute a genuine quantum version of the algorithm. The motivation for leveraging quantum computers for PCG, which is called quantum procedural content generation (QPCG) in the following, lies in their potential to introduce new levels of complexity, creativity, and efficiency. Because quantum physics exhibits intrinsic randomness [5], quantum random number generators (QRNGs) can be viewed as sources of true randomness [6], making them a natural choice for generating randomized content. In addition, quantum superposition allows the random generation of complex patterns and variations that classical algorithms might find challenging to generate efficiently.
Currently available noisy intermediate-scale quantum (NISQ) hardware has very limited capabilities and is subject to significant noise and uncertainty [7]. While these kind of disruptive effects are typically to be avoided, they can also lead to novel and unexpected results for the generated content, potentially fostering creativity. In this sense, despite the technological limitations, it seems promising to use quantum computers already today for simple PCG tasks and to study the results. For more complex tasks, hybrid quantum-classical algorithms are a promising strategy to overcome the limitations, for example by integrating quantum subroutines into classical PCG methods to increase variability and complexity.
QPCG is therefore a promising approach, currently mainly for research, but in the future possibly also for practical applications. However, a general discussion of QPCG is beyond the scope of this paper. Instead, the focus will specifically be on the WFC algorithm and how it can be modified to enable the use of gate-based quantum computers [8]. This paper has four main contributions:
@@
  • A probabilistic formulation of the simple tiled model of WFC is presented as a foundation for a quantum version. We denote this formulation as classical wave function collapse (CWFC) to clearly distinguish between classical and non-classical approaches.
  • The quantum wave function collapse (QWFC) method is proposed, which founds on the encoding of a probability distribution in a quantum circuit.
  • The hybrid quantum-classical wave function collapse (HWFC) method is proposed to take into account the limitations of NISQ hardware.
  • The proposed methods are tested on simulators and IBM Quantum devices.
@@
The remainder of the manuscript is organized as follows. After a brief summary of related work, a formal description of CWFC is given in an appropriate probabilistic form, which subsequently allows the development of QWFC and HWFC. The proposed methods are then demonstrated in practice. Finally, the paper ends with a conclusion.
RELATED WORK
There are many approaches to PCG, such as machine learning [9], evolutionary algorithms [10] and search-based methods [11]. WFC was originally proposed by Gumin [1] as an example of the latter. Subsequent research has extended the basic WFC algorithm to address its limitations and expand its capabilities [2, 3]. For example, by eliminating the need for two-dimensional grids [12], adding design constraints [13], or allowing interactive user control over the results [14, 15]. Better scaling and runtime can be achieved by a hierarchical approach [16], a nested approach [4] or by using bitwise operations [17].
In contrast to classical PCG, QPCG is much less common in the literature. Previous work on QPCG was mainly focusing on the quantum generalization of a blurring process [18, 19] and map generation using a quantum-enhanced decision making process [20]. Moreover, related projects have been carried out in connection with the development of quantum-inspired games [21].
CLASSICAL METHODS
This section provides a formal framework for the following methods, referred to as probabilistic iterative PCG (PIPCG). Subsequently, it is shown how the common CWFC is a special case of PIPCG. Only the main features of PIPCG and CWFC are summarized, the technical details can be found in the appendix.
Probabilistic Iterative PCG (PIPCG)
@@
where each segment $(i,v_{i})$ is defined by a unique identifier $i\in[1,N]$ and a value $v_{i}\in A$ from an alphabet $A:=[1,W]$ consisting of $W$ symbols with the notation $[i,j]:=\{i,\dots,j\}$. The set of all possible content instances $\mathcal{C}$ contains $W^{N}$ elements. For example, if the content represents an image, its individual pixels can constitute the segments. Each pixel can attain one color from the available color space of the image (as its alphabet).
Iterative Process.
@@
    Randomly select a new segment identifier $s(k)\in[1,N]$ from a set of available identifiers with a user-defined probability distribution.
@@
    Randomly select a corresponding value $v_{s(k)}\in A$ from a set of available values with a user-defined probability distribution.
@@
By repeating these steps, the content instance is assembled in the sense of $C(0)\overset{k=1}{\rightarrow}C(1)\overset{k=2}{\rightarrow}\cdots\overset{k=N}{\rightarrow}C(N)\equiv C$, where $C(k):=\{(s(1),v_{s(1)}),\dots,(s(k),v_{s(k)})\}$ represents the generated partial content instance at the end of iteration $k$. After $N$ iterations, the newly generated content instance is complete.
Probabilistic Generation.
An essential concept of PCG is that only a limited subset of all possible content instances $C\in\mathcal{C}$ is generated according to the designer’s aesthetic preferences, a given set of rules, or some other underlying logic; otherwise, the instances could be drawn directly from $\mathcal{C}$ with much less effort. For PIPCG, the user-defined probability distributions for the segment identifiers and the corresponding values represent exactly this underlying logic.
Generating content with PIPCG effectively corresponds to drawing a sample from the random variable $\boldsymbol{C}\sim p(C)$, where
$p(C)=p(v_{1},\dots,v_{N})$ (2)
denotes the probability to generate the content $C$ in the form of Equation 1 with values $v_{1},\dots,v_{N}$. Hence, $\underline{\mathcal{C}}:=\{C\,|\,C\in\mathcal{C}\,\land\,p(C)>0\}$ represents the set of all content instances that can potentially be generated out of all possible content instances $\mathcal{C}$.
Classical Wave Function Collapse (CWFC)
CWFC can be understood as a special case of PIPCG. In this manuscript, only a simplified form of the simple tiled model of CWFC is considered (with the minor extension of graph-based adjacency [12]), as it is sufficient to capture the key components of the approach. To realize CWFC, the two randomized selection steps of PIPCG are chosen as follows:
  1. 1.
    For the segment identifier selection, the segments with the smallest Shannon entropy with respect to the available value options are chosen.
  2. 2.
    Value selection is determined by a pattern-based set of user-defined rules.
Adjacency.
For the pattern-based value selection, the segments are organized in such a way that they have a predefined adjacency relationship to each other that does not change. For example, consider that the pixels of an image constitute the segments. Choosing adjacency between the nearest neighbors leads to an adjacency configuration with four directions (right, up, left, down) as shown in Figure 1.
Formally, each direction within a adjacency configuration is specified by an index $d\in[1,D]$, where $D$ denotes the total number of directions. The direction-based adjacency relationship between different segments is then defined by the coefficients $\alpha^{d}_{i,j}\in\{0,1\}$ for $i,j\in[1,N]$ and $d\in[1,D]$. If $\alpha^{d}_{i,j}=1$, segment $i$ is connected to segment $j$ in direction $d$; otherwise not. For each direction $d$, these coefficients form an adjacency matrix $\alpha^{d}\in\{0,1\}^{N\times N}$ of a directed graph with $N$ vertices that represent the segments. Hence, the segment adjacency is in fact an abstract concept that is not necessarily related to the visual form of the content.
Patterns.
Patterns define how values are chosen based on the already generated content. A pattern
$P:=\{(d_{1},v_{1}),\dots,(d_{n},v_{n})\}$ (3)
consists of a set of $n\leq D$ direction-value pairs with $d_{i}\in[1,D]$ and $v_{i}\in A$ for $i\in[1,n]$, which represents a layout of segments. A pattern-based rule
$r_{i}^{\mathrm{P}}(v,P,u)=(v,u,P)$ (4)
is defined by a value $v\in A$, a factor $u\in\mathbb{R}_{>0}$, and a pattern $P$. It describes the weight $u$ of selecting a value $v$ for the target segment given the layout of adjacent segments $P$. The higher the value $u$, the more probable becomes the selection of the corresponding value. A selection mechanism with $m$ rules is then fully defined by a ruleset
$R:=\{r_{1},\dots,r_{m}\}$ (5)
comprised of rules from Equation 4 with $i\in[1,m]$.
CWFC Example: Checkerboard
As a simple example, consider the generation of images with black and white checkerboard patterns. The pixels of an image constitute the segments and the alphabet consists of only $W=2$ symbols standing for the two colors. As an adjacency configuration, the nearest-neighbor adjacency from Figure 1 is presumed.
A checkerboard image can be achieved with only two patterns of the form of Equation 3, namely $P_{1}:=\{(1,1),(2,1),(3,1),(4,1)\}$ and $P_{2}:=\{(1,2),(2,2),(3,2),(4,2)\}$. The corresponding ruleset, Equation 5, reads $R=\{r_{1}^{\mathrm{P}},r_{2}^{\mathrm{P}}\}$ and therefore contains the $m=2$ rules $r_{1}^{\mathrm{P}}(v=2,P=P_{1},u=1)$ and $r_{2}^{\mathrm{P}}(v=1,P=P_{1},u=1)$, see Figure 2.
Figure 2: Visualization of pattern-based rules. (a) Visualization scheme. The center tile corresponds to the target segment $i$ with a target value of $v$, whereas the adjacent tiles represent the required values $v_{1},\dots,v_{4}\in\{1,2\}$ of the four adjacent segments in the respective directions $d=1,\dots,d=4$. (b) Pattern-based rules $r_{1}^{\mathrm{P}}(v=2,P=P_{1},u=1)$ and $r_{2}^{\mathrm{P}}(v=1,P=P_{2},u=1)$, Equation 4, that can be used for the generation of images with black and white checkerboard patterns.
The iterative process of CWFC is sketched in Figure 3 for a $3\times 3$ image (represented by $N=9$ segments). Only two content instances can be generated, $C_{1}:=\{(1,1),\dots\}$ and $C_{2}:=\{(1,2),\dots\}$ with equal probability $p(C_{1})=p(C_{2})=\frac{1}{2}$, Equation 2. The choice of the first value $v_{s(1)}$ already determines the choices of all other values.
Figure 3: CWFC for the generation of $3\times 3$ images with checkerboard patterns defined by the rules from Figure 2. In each iteration $k\in[1,9]$, three steps take place: (i) a segment identifier $s(k)$ with the smallest entropy is drawn, (ii) a corresponding value $v_{s(k)}$ is drawn according to the pattern-based ruleset, and (iii) the newly generated identifier-value pair is added as a segment to the content instance $C(k)$. For $s(k)$ and $v(k)$, the uniformly distributed set of possible choices are listed. The choices for $s(k)$ depend on the Shannon entropy $H_{k}$ of each undefined segment (see appendix), as shown on the right. The choices for $v(k)$ depend on the fulfillment of the two rules $r_{1}^{\mathrm{P}}$ and $r_{2}^{\mathrm{P}}$, as listed.
QUANTUM METHODS
In the previous section, only classical PCG was considered. This section is about QPCG. First, QWFC is proposed as a special case of PIPCG. For this purpose, the probability distribution of the content instances $p(C)$, Equation 2, is realized with the help of a quantum circuit and the sampling is performed by exploiting the intrinsic randomness of quantum measurements. That is, the actual quantum mechanical wave function collapse is used to implement a quantum version of the Wave Function Collapse algorithm.
Subsequently, this section proposes HWFC as a quantum-classical hybrid algorithm based on content partitioning to overcome the limitations of NISQ devices. Only the main features of QWFC and HWFC are summarized, the technical details can be found in the appendix. A software implementation is available online [22].
Quantum Wave Function Collapse (QWFC)
To realize QWFC, the two randomized selection steps of PIPCG are chosen as follows:
  1. 1.
    The segment identifier selection is performed with a predefined order given by the vector $\sigma\in S_{N}$, where $S_{N}$ is the set of all permutations of $[1,N]$.
  2. 2.
    The value selection is based on a pattern-based ruleset in the same way as for CWFC.
Value Encoding.
The value for each segment is encoded by a set of qubits, which requires $q:=\lceil\log_{2}W\rceil$ qubits for each segment $i\in[1,N]$. The group of $q$ qubits that represent the value for segment $i$ is denoted by $\mathcal{Q}_{i}:=\{q_{i}^{1},\dots,q_{i}^{q}\}$, where $q_{i}^{j}$ references the $j$th qubit of this group for $j\in[1,q]$. The state of these qubits encodes the value $v_{i}$ of the segment $i$ in binary as the tensor product sequence of qubit states $\bigotimes_{j=1}^{q}\ket{b_{j}}=\ket{v_{i}-1}_{i}\in\mathcal{H}_{i}$ via $v_{i}-1=\sum_{j=1}^{q}b_{j}2^{j-1}$ with $b_{j}\in\{0,1\}$. Here, $\mathcal{H}_{i}$ denotes the joint Hilbert space of all qubit from $\mathcal{Q}_{i}$ with $v_{i}\in A$ for $i\in[1,N]$.
In total, $Q:=Nq$ qubits are required to represent the value of all $N$ segments. The joint Hilbert space of all qubits from $\bigcup_{i=1}^{N}\mathcal{Q}_{i}$ is given by $\mathcal{H}:=\bigotimes_{i=1}^{N}\mathcal{H}_{i}$.
Probability Encoding.
Presume a state |$\ket{\Psi}\in\mathcal{H}$ that can realize samples from $p(C)$, Equation 2, with
$p(C\!=\!\{(1,v_{1}),\dots,(N,v_{N})\})=\big{|}\bra{\Psi}\bigotimes_{i=1}^{N}\ket{v_{i}-1}_{i}\big{|}^{2}.$ (6)
To prepare |$\ket{\Psi}$ from a ground state |$\ket{0}:=\bigotimes_{i=1}^{N}\ket{0}_{i}$ based on a pattern-based ruleset $R$, Equation 5, a series of operators is applied, one for each iteration $k\in[1,N]$ such that |$\ket{\Psi}:=U_{N}\cdots U_{1}\ket{0}$ represents the final state after $k$ iterations. In each iteration $k$, the state of the qubits $\mathcal{Q}_{\sigma_{k}}$ (that represent the value for segment $\sigma_{k}$) is prepared conditioned on the state of (a subset of) the qubits $\mathcal{Q}_{\sigma_{1}},\dots,\mathcal{Q}_{\sigma_{k-1}}$ (that represent the values of the already prepared segments $\sigma_{1},\dots,\sigma_{k-1}$), which leads to an entangled joint state as sketched in Figure 4(a).
Figure 4: QWFC circuit with $\sigma=(1,\dots,N)$. (a) Circuit layout with an exemplary alphabet $A$ with $W=16$ symbols, which requires a set $\mathcal{Q}_{i}$ of $q=4$ qubits for each segment with $i\in[1,N]$. In each iteration $k\in[1,N]$, the operator $U_{k}$ prepares the qubits from $\mathcal{Q}_{k}$ in a superposition state |$\ket{C^{\prime}}_{k}$, entangled with (a subset of) the qubits from $\mathcal{Q}_{1},\dots,\mathcal{Q}_{k-1}$. All qubits are measured to obtain $p(C)$ (see appendix). (b) Gate decomposition of $U_{k}$ into the components of $U_{k}^{\mathrm{coin}}$, which consists of pairs $U_{k}^{\mathrm{ctrl}}(C^{\prime})\otimes U_{k}^{\mathrm{init}}(C^{\prime})$. Each of these pairs represent a conditional loading of a probability distribution. The combined control operators act on the qubits from $\mathcal{Q}_{k}^{+}$, whereas the initialization operators act on the qubits from $\mathcal{Q}_{k}$ (see appendix).
The required number of qubits for the proposed QWFC method increases linearly with the number of segments and logarithmically with the number of symbols in the alphabet. At the same time, the respective circuits may become exponentially deeper because of the costly conditional preparation of probability distributions [23].
QWFC Example: Checkerboard
In the following, the above example for the generation of $3\times 3$ checkerboard images is considered for QWFC, where the predefined segment order $\sigma=(1,2,3,6,5,4,7,8,9)$ is used as visualized in Figure 5(a). The resulting CWFC circuit is shown in Figure 5(b) and the resulting probability distribution of content instances $p(C)$, Equation 6, in Figure 5(c). A (noise-free) measurement of the circuit corresponds to drawing a sample from this distribution.
Figure 5: QWFC for the generation of $3\times 3$ images with black and white checkerboard patterns in analogy to Figure 3. (a) Predefined segment order $\sigma=(1,2,3,6,5,4,7,8,9)$. (b) Circuit layout with the same symbols as in Figure 4. (c) Generated content instances $C_{1}$ and $C_{2}$ with $p(C_{1})=p(C_{2})=\frac{1}{2}$.
Hybrid Quantum-Classical WFC (HWFC)
[…]
DEMONSTRATION
In the present section, five simple content creation use cases are presented to demonstrate the proposed methods. For QWFC and HWFC, both idealized (i.e., noise-free) simulators and IBM Quantum devices were used, the latter being accessed via the IBM Quantum Cloud Services [24] during December 2023.
Checkerboard Revisited
The first use case is the previously presented checkerboard example, cf. Figures 2 and 5. Again, $3\times 3$ instances are considered, for which a comparison between QWFC and HWFC is performed on the IBM Quantum device ibm_kyoto (Eagle r3 processor version 1.2.4, 127 qubits). For QWFC, 10 0001000010\,00010 000 shots are measured, each yielding a content instance. For HWFC, $H=3$ equal-sized partitions are used in analogy to Figure 6 and the method is repeated $131$ times, each time yielding a content instance. The results are shown in Figure 7.
[…]
In total, there are $5122^{9}=$ 512 possible content instances $C\in\mathcal{C}$, but only two of them, $C_{1}$ and $C_{2}$, are valid. The remaining $510$ content instances are invalid because they violate the prescribed patterns. In Figure 7(a), QWFC is able to generate the two valid instances, but at the same time also produces a lot of invalid instances due to hardware imperfections. From Figure 7(b) it can be seen that the results from HWFC are much closer to the idealized solution that would contain only $C_{1}$ and $C_{2}$ with equal probability. It is a consequence of the reduced circuit size from this method that invalid solutions occur with much less frequency than for QWFC.
Pipes
The second use case is to create an image from tiles of seven different pipe (or line) sections as well as one blank tile. That is, the alphabet consists of $W=8$ symbols and each tile in the image constitutes a segment. The alphabet is shown in Figure 8(a). As adjacency configuration, the nearest-neighbor adjacency from Figure 1 is presumed with $D=4$. The pattern-based ruleset is chosen in such a way that neighboring tiles have to form a connecting network of pipes. Three example rules for this purpose are shown in Figure 8(b). In total, $m=2048$ rules are required to achieve this goal.
Figure 8: Pipe generation. The generated image is required to show a network of connected pipes. (a) Alphabet consisting of $W=8$ symbols, each of which represents a tile with a pipe section and an empty tile. (b) Example of three pattern-based rules $r_{1}^{\mathrm{p}}$, $r_{2}^{\mathrm{p}}$, and $r_{3}^{\mathrm{p}}$, cf. Figure 2. (c) Example $10\times 4$ tile images that were generated using HWFC with $H=10$ on the IBM Quantum device ibm_kolkata and a simulator; also shown are images from CWFC.
Exemplarily, images consisting of $10\times 4$ tiles ($N=40$) are considered, for which a comparison between HWFC on the IBM Quantum device ibm_kolkata (Falcon r5.11 processor version 1.14.8, 27 qubits), HWFC on a simulator and CWFC is performed. For HWFC, $H=10$ equal-sized partitions are used, each corresponding to one column of the image. A few resulting instances are shown in Figure 8(c).
Hexagon Map
The third use case addresses images based on hexagonal tiles that can be interpreted as a map consisting of different terrains. For this purpose, an alphabet of $W=4$ symbols is chosen, each corresponding to a unicolored tile (blue, yellow, green, and gray) and each tile constitutes a segment. As adjacency configuration, a nearest-neighbor adjacency is presumed ($D=6$). The prescribed pattern-based ruleset only allows connections of the form blue-yellow-green-gray, which can be fulfilled with $m=1586$ rules, see Figure 9(a). To promote a higher occurrence of blue tiles, the factors in Equation 4 are chosen as $u=5$ for blue tiles and $u=1$ otherwise.
As a first example, images consisting of $N=7$ tiles are considered, for which QWFC on a simulator and CWFC are performed. The resulting distributions of content instances $p(C)$ in Figure 9(b) is exactly calculated for QWFC, Equation 6, and based on 10 0001000010\,00010 000 samples for CWFC, Equation 2. Both distributions are significantly different, which is no surprise since the two approaches use a different segment identifier selection. The second example considers images from $N=127$ tiles, for which HWFC with $H=20$ is performed on a simulator. Example instances are shown in Figure 9(c).
Platformer
The fourth use case is motivated by computer game level design. The goal is to create a platformer-type level from a set of eight tiles [25], each corresponding to a different game element (ground, grass, mushroom, block, air, and a tree that is composed of three tiles). Consequently, the alphabet consists of $W=8$ symbols, see Figure 10(a). As adjacency configuration, the two tiles above and below are taken into account as neighbors ($D=2$). A pattern-based ruleset is prescribed that meets five requirements: (i) The bottom row consists of ground tiles and ground tiles can only be placed on top of each other, (ii) only a grass tile or a mushroom tile can be placed above a ground tile, (iii) tree tiles must be placed in order with the lowest tile above a grass tile, (iv) air tiles can only be placed above grass, mushroom or other air tiles, (v) block tiles can only be placed between two air tiles. These requirements can be fulfilled with $m=15$ rules, see Figure 10(b). The first requirement can be resolved with a functional factor $u(i)$ in Equation 4 that vanishes for all non-ground tiles in the bottom row (see appendix). For all non-vanishing cases, $u=$0.1$$ for $r_{6}^{\mathrm{p}}$ and $u=1$ otherwise to reduce the occurrence of block tiles. The segment order $\sigma$ is chosen such that the levels are built up from bottom to top.
Figure 10: Level generation. (a) Alphabet consisting of $W=8$ symbols, each of which represents a tile [25]. (b) Pattern-based ruleset consisting of $m=15$ rules. (c) Example $10\times 10$ tile images that were generated using HWFC with $H=20$ on a simulator.
Exemplarily, levels with $10\times 10$ tiles are considered, for which HWFC with $H=20$ equal-sized partitions (each corresponding to a half row of the level) is performed on a simulator. Four resulting instances are shown in Figure 10(c).
Voxel Skyline
The final use case considers the creation of a three-dimensional voxel graphic with a binary alphabet of $W=2$ symbols, which represent the presence or absence of a voxel. Each voxel constitutes a segment, as adjacency configuration the voxel above and below are taken into account ($D=2$). The chosen pattern-based ruleset with $m=4$ ensures that voxels are built from the ground up, see Figure 11(a). This leads to a skyline-like structure.
Figure 11: Skyline generation. (a) Pattern-based ruleset consisting of $m=4$ rules in an analogous representation as in the previous figures. (b) Example images from a $4\times 4\times 4$ voxel grid that were generated using QWFC on the IBM Quantum device ibm_osaka and a simulator. (c) Example images from a $10\times 10\times 5$ voxel grid that were generated using QWFC on a simulator.
Exemplarily, two cases are considered: First, images from a $4\times 4\times 4$ voxel grid, for which QWFC is performed both on a the IBM Quantum device ibm_osaka (Eagle r3 processor version 1.0.3, 127 qubits) and a simulator, see Figure 11(b). An error violating $r_{3}^{\mathrm{p}}$ can be seen in the bottom left image that results from hardware imperfections. Second, images from a $10\times 10\times 5$ voxel grid, for which QWFC is performed on a simulator, see Figure 11(c).
CONCLUSION
In this work, QWFC was proposed as a method to realize WFC on a gate-based quantum computer as an example of QPCG. The idea is to construct a quantum circuit from which valid content instances (that fulfill the prescribed patterns) can be sampled via measurements. In other words, the quantum circuit acts as a special-purpose QRNG for content of a desired form that makes use of the intrinsic randomness of quantum physics. This means that the actual quantum mechanical collapse of the wave function is used for the implementation, from which the classical method derived its name.
The proposed method is not without its challenges. The generating quantum circuit becomes larger from both the content and alphabet size and at the same time deeper from correlations within the content instance distribution, which makes it difficult to evaluate practical use cases on current hardware. For this reason, HWFC was proposed as a hybrid method that partitions the generation into smaller tasks for QWFC.
The experimental results have shown the limitations of QWFC and HWFC, but have also proven that at least simple examples can already be implemented on today’s quantum hardware. The biggest bottleneck of the circuit complexity is the conditional preparation of probability distributions. Investigating possible improvements in this area with regard to specific QPCG use cases could therefore be a potential direction for future research.
The conceptual advantage of QWFC is that a circuit, once designed, represents the distribution of content instances such that new instances can be generated by simple measurements without any additional algorithmic effort. A shortcoming of QWFC is that only a fixed order of segments is considered instead of a dynamic (e.g., entropy-based) order that depends on the previously selected segments, which is a chance for future improvements.
In conclusion, the paper aims to provide an initial approach to QPCG that demonstrates its feasibility but leaves room for improvement and raises the question of useful application areas given the limits of NISQ devices.
APPENDIX
This appendix section contains the technical details of the presented PCG methods.
Technical Details of PIPCG